OpenStudy (anonymous):

Find two vectors parallel A and perpendicular A that add up to A such that one is parallel and one is perpendicular to B. Let A= i-3j+2k and B=2i-4k. So since A= parallel A + perpendicular A, I worked out parallel A to be { -1, 0, 2} and perpendicular A to be {0, -3, 0}. But I don't understand how to relate that back to B. The answers are (-3/5)i + (6/5)k and (8/5)i + (-3)j +(4/5)k

OpenStudy (amistre64):

i cant make out the instructions that well

OpenStudy (amistre64):

if parallel A is over A and thru A under over with a dash of per A on the sid eof A then we want to A the addition of A and blah blah blah .... who came up with this mess?

OpenStudy (anonymous):

http://math.scu.edu/~gmohler/practicemidterm2_math13.pdf it's problem #3 to this worksheet

OpenStudy (amistre64):

burn your worksheet lol

OpenStudy (amistre64):

does parallel and perp have to be parallel and perp to A? or is that just what they wanted to go as a naming convention?

OpenStudy (anonymous):

I'm assuming it's just a naming convention

OpenStudy (amistre64):

its a terrible convention then; but otherwise how would you get 2 vectors that add up to A otherwise right

OpenStudy (anonymous):

I found parallel A by using projB A = (AB/ BB) B

OpenStudy (amistre64):

perp would be B.A

OpenStudy (amistre64):

or at least a vector named perpA i think

OpenStudy (anonymous):

ah, I'll try that and see

OpenStudy (amistre64):

perpA.B = 0 if that makes sense

OpenStudy (amistre64):

subtract your //A from A and see if it pers with B by chance

OpenStudy (anonymous):

that would give me perpendicular A

OpenStudy (amistre64):

it might

OpenStudy (amistre64):

if i see it right, and thats a big if, i get something akin to \[n<2,0,-4>\ + <2,0,-4>*<a,b,c>=<1,-3,2>\]

OpenStudy (amistre64):

to me this means that: \[n<2,0,-4>\ + 0\ =<1,-3,2>\] \[2n=1;\ 0=-3;\ -4n=2\] that doesnt seem right does it

OpenStudy (amistre64):

i see what i did, i dotted in the wrong spot lol

OpenStudy (amistre64):

<a,b,c> is spose to add to n<2,0,-4> and get <1,-3,2> <a,b,c> is spose to dot to <2,0,-4> and equal 0

OpenStudy (anonymous):

I found in my notes that perpendicular A + //A = A; and perpendicular A x B = 0. But how would the latter help me?

OpenStudy (amistre64):

prove to me you found //A, show me your steps so i can chk it

OpenStudy (amistre64):

LA and //A are good for symbols

OpenStudy (anonymous):

//A = ProjB A = (2+0-8)/(4+8) x {2, 0, -4} = -1/2 x {2, 0, -4} = {-1, 0, 2}

OpenStudy (anonymous):

and then I did LA = A - //A so LA = {1, -3, 2} - {-1, 0, 2} = {0, -3, 0}

OpenStudy (amistre64):

i did this: \[nB = //A =<2n,0,-4n>\] \[//A+<a,b,c>=A<1,-3,2>\] \[\begin{array}l 2n+a=1\ a=1-2n\\ 0n+b=-3;\ b=-3\\ -4n+c=2;\ c=2+4n \end{array}\] <1-2n , -3, 2+4n> < 2 , 0 , -4 > ------------------- 2-4n +0 -8-16n = 0 -6 -20n = 0 -20n=6 n= -6/20 = -3/10

OpenStudy (anonymous):

oh wait, I made a mistake

OpenStudy (amistre64):

(-3/10) <2,0,-4> = <-3/5 , 0 , 6/5> which is what your answer shows

OpenStudy (amistre64):

thats //A

OpenStudy (amistre64):

A - //A = LA, it has to

OpenStudy (anonymous):

8/5i -3j+4/5k

OpenStudy (anonymous):

thank you so much! :D

OpenStudy (amistre64):

your welcome, i hope it made sense :)